### v1

# Predicting Chemically Accurate Adsorption Energy Using an Interpretable Deep Learning Model Pretrained by GGA Calculation Data

Abstract
Molecular adsorption energy is a critical descriptor for high-throughput screening of heterogeneous catalysts and electrode materials. However, precise experimental adsorption energies are scarce due to the complexity of experiments, while density functional theory (DFT) calculations remain computationally expensive for large-scale material screening. Machine learning models trained on DFT data have emerged as a promising alternative, but face challenges such as functional dependency, limited labeled high-fidelity data, and interpretability issues. Herein, we present DOS Transformer for Adsorption (DOTA), a novel deep learning model that leverages local density of states (LDOS) as input to predict chemically accurate adsorption energies across metallic and intermetallic surfaces. DOTA integrates multi-head self-attention mechanisms with LDOS feature engineering to capture orbital interaction patterns, achieving superior accuracy and transferability. Pretrained on PBE-level DFT data, DOTA can be fine-tuned using minimal high-fidelity experimental or hybrid functional data to predict adsorption energies with chemical accuracy. The model resolves long-standing challenges, such as the "CO puzzle" and outperforms traditional theories, like the d-band center and Fermi softness models. It provides a robust framework for efficient catalyst and electrode screening, bridging the gap between computational and experimental approaches.
Introduction
Molecular adsorption energy has been widely used as a descriptor for high-throughput screening of heterogeneous catalysts and electrode materials. [1, 2, 3, 4, 5, 6, 7, 8] Due to the substantial inherent uncertainties arising from the facts, such as side reactions, surface defects, and impurities in most experiments, precise experimental adsorption energies could only be obtained through single-crystal adsorption microcalorimetry with an elaborate experimental design so far, resulting in scarce benchmark experimental data. [9, 10, 11, 12, 13, 14, 15, 16, 17] This limits the application in screening candidate materials. Consequently, the adsorption energies derived from density functional theory (DFT) calculations have been extensively applied for materials screening. [18, 19, 20, 21, 22] Yet, it is not easy to afford the huge computational cost of DFT calculations to achieve a great number of adsorption energies at various active site ensembles in multi-component alloys. Therefore, machine learning models trained using data labeled by DFT calculations have recently emerged as a promising method to predict the adsorption energy for material screening. [23, 24, 25, 26, 27, 28, 29, 30, 31] Furthermore, the deep learning model demonstrated superior generalization and transferability for a broader range of adsorbates and catalysts compared with surrogate models utilizing a small dataset for specific adsorbates and materials. [32, 33, 34, 35, 36] Notably, machine learning interatomic potentials, which obtain optimized adsorbate structures and adsorption energies using structure coordinates as inputs, have been established as a critical technique in recent years. [37, 38, 39, 40, 41]
However, the prediction of chemically accurate adsorption energies using labeled DFT calculation data remains a significant challenge. For one thing, the reduced symmetry of surface adsorption structures increases the computational cost of DFT calculations to label each data point compared to bulk materials and molecules. For another, the low symmetry necessitates more than two orders of magnitude more surface adsorption data for machine learning model training than bulk materials and molecules. [42] Consequently, this results in significantly scarcer labeled adsorption data compared to molecular and bulk material data, despite the development of several databases such as Open Catalyst and Catalysis-Hub. [43, 44, 45] Furthermore, it is challenging to leverage limited labeled adsorption data for model training. The labeled DFT data frequently originate from different exchange-correlation functionals and calculation parameter settings, such as different basis sets, k-point samplings, and versions of pseudopotentials, indicating inconsistent and discontinuous potential energy surface data for the training of machine learning interatomic potential. As a result, only a small portion of the labeled data with consistent exchange-correlation functional and calculation parameter settings in the databases were utilized in most of the previous work in a single task. [46, 47, 48, 49, 50]
More importantly, the accuracy of the predicted adsorption energy based on the deep learning model is subject to the exchange-correlation functional utilized for the training set, rendering it functional-dependent. The constraint of computational cost renders the adsorption databases to be predominantly composed of the results based on generalized gradient approximation (GGA) exchange-correlation functionals up to now, particularly the PBE [51] functional. [52, 53] However, the PBE functional exhibits limitations in terms of accuracy, exemplified by the classic "CO puzzle" problem. [54, 55, 56, 57, 58] Moreover, the meta-GGA functionals of SCAN [59] and r2SCAN [60], which have been regarded as the promising functionals to establish a reliable adsorption database recently, even overbind slightly compared to PBE and mistakenly predict the preferential site for CO adsorption as well. [61] The substantial computational cost associated with high-level quantum chemistry methods, such as random phase approximation (RPA) [62], which can yield chemically accurate adsorption energies using the embedding strategy, [63, 64, 65, 66] precludes their use in constructing a sufficiently large dataset for model training. The -machine learning method presents an alternative approach for predicting chemically accurate adsorption energies for specific systems by incorporating RPA calculation results. Nevertheless, approximately one hundred RPA results are still required to train a model to predict CO adsorption energies on a specific metal surface using the -machine learning method. [67, 68, 69, 70] Consequently, the extension of the method to vast chemical compound spaces is hindered by significant computational expenditures. Thus, the application of this method to catalyst and electrode screening is severely limited due to these constraints. It is imperative to develop a new model and protocol to predict chemically accurate adsorption energies, making full use of the limited GGA-level data available.
To this end, breaking the limitation of functional-dependency is a prerequisite for a new machine learning model for precise adsorption prediction. As illustrated in Fig. 1a, the machine learning interatomic potential, which corresponds to the mapping from input atomic coordinates to output energies, must be functional-dependent, arising from the intrinsic functional-independency of atomic coordinates, while the functional-dependency of energies. This indicates that a functional-independent deep learning model may be achieved by employing functional-dependent physical quantities as both model input and output. Notably, both adsorption energies and electronic structure characteristics exhibit strong functional-dependency. The density of states (DOS) serves as an information-dense descriptor that compresses a substantial amount of spatially distributed electronic structure information into a computationally affordable one-dimensional energy space, making it a promising candidate as a model input to overcome the limitation of functional-dependency.
Interestingly, DOS is one of the most frequently used physical quantities that help to understand the chemisorption strength in heterogeneous catalysis. In particular, the d-band center theory, [71, 72], which is derived from the approximation of the Newns-Anderson model, [73, 74, 75] is prevalent in the field of heterogeneous catalysis. Recently, Xin’s group proposed a Bayesian learning method to emphasize the significance of the orbital orthogonalization term in the Newns-Anderson model for noble metal surfaces, which was previously omitted in the d-band center theory. [76] Fermi softness offers an alternative explanation for the orbital contribution to adsorption by extending the frontier molecular orbital theory to the solid surface. [77] This approach underscores the critical role of the DOS near the Fermi levels. Fung et al. proposed a deep convolutional neural network (DOSNet) to predict the adsorption energy at the GGA level by utilizing the local density of states (LDOS) of the adsorbate and the binding sites of the substrate, mainly based on the adsorbed structures. [78] Nevertheless, the CNN featurizer in DOSNet suffers from low explainability and long-range interaction neglect [79, 80, 81, 82, 83], which limits the comprehensive understanding of the contribution of different energy levels to the adsorption.
Despite extensive studies on the relationship between adsorption energies and DOS, the fundamental challenge of functional-dependency for adsorption strength remains unaddressed. Herein, we present a new deep learning model, called DOS Transformer for Adsorption (DOTA), to efficiently predict accurate adsorption energies across metallic and intermetallic surfaces. The DOTA architecture synergistically integrates the interpretable multi-head self-attention mechanism of the Transformers framework with LDOS feature engineering to first capture the underlying modes of orbital interactions through pretraining using PBE-level LDOS profiles and adsorption energies without prior physical constraints. The functional-independent orbital interaction modes enable transfer learning across quantum chemistry methods and experimental data. Therefore, the chemically accurate adsorption energies of a specific adsorbate across various metallic and intermetallic surfaces could be predicted by fine-tuning the PBE-level model through incorporating the LDOS profile of the gaseous adsorbate at the HSE06 [84] level and few-shot precise experimental adsorption energies. We also comprehensively revisited the d-band center theory and Fermi softness through our interpretable attention paradigm and deciphered the CO/Pt(111) puzzle.
Results
Workflow of DOTA
The DOTA framework employs a two-stage training protocol comprising pretraining and fine-tuning steps to predict chemically precise adsorption energies, as schematically illustrated in Fig. 1b. During the pretraining step, all LDOS inputs and corresponding chemisorption energy outputs were derived from PBE results, thereby constructing a DOTA-PBE model. This model establishes fundamental orbital interaction patterns between adsorbates and metallic surfaces and predicts precise PBE adsorption energies of the trained adsorbates across metallic and intermetallic surfaces. Subsequent fine-tuning integrated GGA-level LDOS-adsorption energy pairs and hybrid data consisting of the HSE06-level DOS profile of the gaseous molecule and few-shot experimental adsorption energies of the studied adsorbate to achieve accurate adsorption energies of the studied adsorbate. To enable efficient catalyst screening without requiring preliminary time-consuming geometry optimization, DOTA strategically processes gaseous adsorbate LDOS and bare surface LDOS instead of adsorbed complex configurations. The model architecture implements quantum mechanically informed feature engineering, allocating 32 angular momentum-projected density of states (PDOS) embedding channels per surface atom and 8 PDOS embedding channels per adsorbate atom. For instance, the bridge-site adsorption configuration corresponds to 64 surface atom channels (2 atoms × 32 channels) and 8 adsorbate atom channels. This channel design ensures compatibility with f-electron systems and spin-polarized calculations. Initial feature processing involved an average pooling layer to normalize energy interval disparities across different DOS calculations. Surface and adsorbate channels were then projected into separate latent spaces through weight-shared fully connected layers with nonlinear activation functions. The multi-head attention mechanism in Transformer architecture enables efficient modelling of long-range orbital interactions across the entire energy spectrum, capturing the interactions between orbitals with different angular momentum quantum numbers. Final energy predictions are generated by merging surface atom features through flattening and sequential fully connected layer processing. Comprehensive implementation details regarding network architecture and hyperparameter selection are provided in the Methods section.
Performance of DOTA-PBE
The pretraining dataset was constructed using 23,861 PBE chemisorption energies involving 11 different adsorbates. These encompassed monoatomic adsorbates (H, C, N, O, and S) and their hydrogenated counterparts (CH, , , NH, OH, , and SH) across 1,982 distinct metallic and intermetallic (111) surfaces comprising 37 metallic elements. The selected adsorbates cover molecules with varying degrees of bonding unsaturation, rendering the chemisorption energies to span from near 0 eV to over -10 eV. This broad energy range provides comprehensive sampling of the diversity of adsorption strengths, ensuring adequate coverage of chemical environments relevant to catalytic processes.
Fig. 2a illustrates the predictive performance of the PBE-level DOTA model via 5-fold cross-validation on the pre-training PBE dataset. The Transformer-based architecture achieves exceptional accuracy of the DOTA-PBE model with a mean absolute error (MAE) of 0.066 eV and a mean absolute percentage error (MAPE) of 1.53 across all the adsorption systems. Adsorbate-specific analysis reveals MAE ranging from 0.042 eV for monoatomic H to 0.096 eV for . MAE differences among different adsorbates arise from the bonding unsaturation and resulting adsorption strength. The stronger adsorption, the higher MAE, but the lower MAPE. This complementary behavior of error metrics exhibits the necessity of taking both absolute (MAE) and relative (MAPE) error measures into account when assessing model performance, particularly given the wide energy range in the dataset. Notably, the reliance on electronic structure features rather than geometric parameters enables effective transferability between adsorbates with analogous bonding characteristics. We also tested the OH adsorption atop (111) surfaces. The tested surfaces were not included in the previous pretraining PBE dataset. As shown in Fig. 2b, the d-band center theory fails at the sites composed of the group 13-14 elements that rely on the s and p orbitals for bonding, while the accuracy of the DOTA-PBE model is element-independent across the metallic and intermetallic surfaces (Fig. 2c). Moreover, it indicates that hydrogenated derivatives may be accurately modelled with minimal additional training data. The electronic structure-based approach of DOTA exhibits robust predictive capability across diverse chemical environments.
We further compared the performance of the DOTA-PBE model with the d-band center and Fermi softness models. We first take the chemisorption of OH over the top site of Ag(111) and Au(111) surfaces as an example, which was not included in our pretraining dataset. As shown in Fig. 3a, the d-band of the Ag(111) surface is mainly concentrated in the energy window far below the Fermi level, implying weakened adsorption if following Fermi softness theory. This is not in good agreement with the PBE calculation result of -1.67 eV for the adsorption of OH atop the Ag(111) surface. Moreover, the decomposition of the contribution of each energy level (Fig. 3b) by the integrated gradients algorithm suggests that those energy levels far below the Fermi level made the main contribution for OH adsorption atop Ag(111) rather than those energy levels around the Fermi level. Interestingly, the lower d-band center of Ag(111) ( = -3.61 eV) than Au(111) surface ( = -3.00 eV), however, results in the stronger adsorption of OH atop Ag(111) than Au(111) (-1.01 eV). These errors may be attributed to the fact that both the d-band center and the Fermi softness theories are based on the presumption of covalent bonding, while the electron transfer from the metallic surface to the singly occupied molecular orbital of OH [85] or Pauli repulsion [86], rather than the orbital overlapping between OH and the Ag(111) or Au(111) surface, dominates the chemisorption of OH atop Ag(111) or Au(111). In contrast, our DOTA-PBE model, which omits any predefined restriction, can accurately predict the adsorption energies of OH atop Ag(111) and Au(111) surfaces, with a prediction of -1.71 eV and -1.13 eV, respectively.
The H adsorption over Pt3Y(111) is a classic example that d-band center theory fails to predict the adsorption site.[77] PBE calculations show that the adsorption energy of H atop the Pt site is -2.80 eV, while it is -0.81 eV atop the Y site, indicating that H tends to preferentially adsorb at the Pt site. However, the Pt site has a lower d-band center ( = -0.32 eV) than the Y site ( = 0.23 eV), indicating stronger adsorption at the Y site if following the d-band center theory. On the contrary, the larger Fermi softness of the Pt site (Sf = 0.92 ) compared with the Y site (Sf = 0.30 ) indicates the stronger adsorption atop the Pt site if following the Fermi softness theory. Hence, the Fermi softness theory can well describe the H adsorption at Pt3Y, while the d-band center theory fails. The DOTA-PBE model can also accurately predict this case, with the predicted adsorption energies atop the Pt and Y sites being -2.75 eV and -0.88 eV, respectively. This surface is not included in our pretraining PBE dataset as well. The contribution of each energy level to the H adsorption was further elucidated by the integrated gradients algorithm. [87] It could be found from Fig. 3c that the Pt valence bands near the Fermi level account for the H chemisorption atop the Pt site, well aligning with the model of Fermi softness. Y conduct bands play a main role in the H chemisorption atop the Y site, which is consistent with the consensus that the strong Lewis acidity of Y could accept the electron of hydrogen. It indicates that the charge transfer induces the Y-H bonding, which explains the reason why the d-band center model is not suitable for describing this system. Hence, the DOTA-PBE model exhibits better generalization for the prediction of adsorption energies across metallic and intermetallic surfaces compared with traditional d-band center theory and Fermi softness theory.
Chemically accurate CO adsorption energy
Building upon the robust performance and transferability of DOTA-PBE, we further fine-tuned the model to achieve chemically accurate predictions of adsorption energies for specific adsorbates across metallic and intermetallic surfaces. The notable scarcity of chemically accurate adsorption energy data requires a model architecture with a minimum requirement of labeled chemically accurate adsorption energy data to fine-tune the model. Taking CO adsorption as an example, we first incorporated 988 LDOS-CO adsorption energy pairs (Table S2) across different sites of 174 metallic and intermetallic surfaces, evaluated by PBE, RPBE [88], and PBEsol [89] functionals, respectively. Data derived from different functionals expanded the perceptive chemical space, thereby enhancing the ability of the model to learn orbital interaction patterns between CO and metallic surfaces. Importantly, the foundation of PBE-level orbital interaction understanding, combined with GGA-level insights into CO-surface interactions, enabled exceptional data efficiency to achieve chemical accuracy predictions with only four high-fidelity data points. Each high-fidelity data point comprised experimentally measured CO adsorption energies (Pd(111), Cu(111), Ir(111), or Ni(111)) through precise microcalorimetry, paired with corresponding PBE-calculated surface LDOS profiles and HSE06-calculated gaseous CO LDOS characteristics (Fig. 4a). We compared adsorption energies of CO over Pt(111) and Rh(111) (Fig. 4c) predicted by the DOTA model with different input functional combinations for surface and adsorbate LDOS profiles with standard PBE, PBEsol, HSE06, RPA, and experimental results. As introduced from Blyholder’s model, [90] CO adsorption is primarily formed through the -donation between CO HOMO (5 occupied orbital) and the metal conduction bands and the d- backbonding between CO LUMO (2 unoccupied orbital) and the metal valence bands. Accordingly, the correct CO adsorption energy and binding sites rely on the correct description of the electronic structures of CO and the metallic surfaces. As reported previously, [91] HSE06 functional is poor at predicting CO adsorption because it fails to describe the delocalized electronic structures of metallic surfaces, while the underestimated HOMO-LUMO gap of CO by PBE and PBEsol functionals [92, 93] should account for the overbinding of CO over Pt(111) and Rh(111) surfaces. [94, 95] Moreover, these deficiencies result in the mistakenly predicted fcc sites for CO chemisorption over these two surfaces. This manifests that the underestimated HOMO-LUMO gap by PBE functional is mainly responsible for the "CO puzzle", which was speculated by previous reports. Interestingly, despite the accurate CO adsorption energy, neither the CO HOMO-LUMO gap nor the metal cohesive energies could be accurately calculated by RPA (Fig. 4), which is consistent with previously reported results. [58] Importantly, HSE06 exhibits the best predictive performance for CO HOMO-LUMO energy levels, while PBE functional excels at the prediction of metal cohesive energy, indicating that HSE06 and PBE functionals are good at the description of the molecular orbitals and metal surfaces, respectively. Consequently, the input of HSE06-calculated gaseous CO LDOS characteristics and PBE-calculated surface LDOS profiles (HSE06/PBE) could provide superior results, with the predicted adsorption energy differing by less than 0.1 eV from the experimental values. More importantly, the model with the HSE06/PBE input combination could correctly predict the top site as the preferential binding site for CO adsorption, aligning with the experimental results. [18] The accuracy of the multimodal integration capitalizes on the advantages of the experimental precision of microcalorimetry in adsorption energy, the reliability of PBE functional for the description of electronic structures of metallic surfaces [96], and the superior performance of HSE06 beyond RPA for CO molecular orbital description [97, 84].
Discussion
The DOTA model integrates LDOSs from both the surface and the molecule as input. Different from the d-band center and Fermi level theories, this approach enables accurate prediction of adsorption energies across diverse adsorbates (MAE = 0.066 eV, MAPE = 1.53) without reliance on predefined theoretical assumptions. Leveraging a Transformer architecture and integrated gradients algorithm, DOTA provides interpretable insights into the contribution of individual energy levels to adsorption strength. Established upon two key physical insights of the origin of the adsorption strength from orbital interactions and the functional-independent DOS-adsorption relationship, DOTA demonstrates its exceptional transferability. This permits integration of experimental sources and DFT data from different functionals and calculation settings. It could further achieve chemical accuracy in predicting molecular adsorption energies across extensive metallic and intermetallic surfaces, primarily trained on GGA-level datasets augmented by few-shot high-fidelity experimental adsorption energy data. Thus, DOTA represents a transformative leap forward in computational catalysis by addressing longstanding challenges such as functional dependency, limited availability of high-fidelity data, and the need for interpretable machine learning models. By synergizing deep learning architectures and physically informed features like LDOS, it achieves unprecedented accuracy and generalizability across chemical environments. Its capacity to unify multi-fidelity data sources and deliver mechanistic insights through interpretable attention paradigms establishes DOTA as a versatile framework for bridging experimental precision with computational efficiency. This could facilitate the construction of a large dataset of the adsorption energies with high fidelity, accelerating the discovery of next-generation catalysts and electrode materials.
Methods
All DFT calculations were performed with Vienna ab initio Simulation Package (VASP) [98, 99] with periodic boundary and the projector augmented wave (PAW) pseudopotentials [98, 100]. The Catalysis Kit (CatKit) software [44] and Atomic Simulation Environment (ASE) Python library [101, 102] were used for generating and classifying the adsorption structures. The chemisorption structures, except for CO adsorption, were obtained in reference [45] for DOTA model pretraining, where PBE functionals were used for both adsorption energy and DOS profile calculations. For CO adsorption, RPBE functional and PBEsol functional were also used for adsorption energy and DOS profile calculations, while the hybrid HSE06 functional was used only for electronic structure calculations. All calculations were conducted with spin polarization, using a kinetic energy cutoff of 450 eV. The p(22) 3-layer slab was employed for each metallic and intermetallic surface. The corresponding Monkhorst-Pack k-point grid [103] was 771, with a first-order Methfessel-Paxton smearing of 0.1 eV. The resolution of calculated lm- and site-DOSs with reference to the Fermi level, which were employed for DOTA training, was 0.01 eV.
The DOTA model was implemented with the PyTorch Python library [104]. The AdamW optimizer with an initial learning rate of 0.0002 and the L1 loss were used for training. An exponential cosine annealing scheduler was implemented to adjust the learning rate, and 150 epochs were used, with 50 epochs for warming up, and a batch size of 256. The contribution of each energy level to the adsorption energy was calculated using the interpretable integrated gradients algorithm implemented in the Captum library [87] with default parameters.
Code availability
Code for DOTA is available as an open source repository on GitHub, https://github.com/zhzhang321/DOTA
Competing interests
The authors declare no competing interests.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (2023YFA1507601 and 2021YFA1500700). We thank the Center for High Performance Computing at Shanghai Jiao Tong University for providing the computing resources of the Siyuan-1 cluster.
Author contributions
Z.Zhang wrote the code and performed theoretical simulations. X.-M.C. designed the study, analyzed the data, and wrote the paper.