### v2

# Minimising the Demand for High-Fidelity Training Data towards Chemically Accurate Adsorption Energy Predictions

Abstract
Adsorption energy is a critical descriptor for high-throughput screening of heterogeneous catalysts and electrode materials. However, precise experimental data are scarce due to the complexity of experiments, while high-fidelity density functional theory (DFT) calculations remain computationally expensive for large-scale material screening. Machine learning models trained on DFT data have emerged as a promising alternative but face challenges such as functional dependency and limited high-fidelity labelled data. Herein, we present DOS Transformer for Adsorption (DOTA), a functional-independent deep learning model established on the map between local density of states (LDOS) and adsorption energy. DOTA integrates multi-head self-attention mechanisms with LDOS feature engineering to capture latent orbital interaction patterns, enabling it to unify multi-fidelity and multi-source data. This minimises the demand for high-fidelity training data. Consequently, the predictive adsorption energy could reach chemical accuracy, requiring less than five high-fidelity experimental adsorption energies for model training. DOTA also resolves long-standing challenges, such as the ”CO puzzle”, and outperforms traditional theories, including the d-band centre and Fermi softness models. It provides a robust framework for efficient catalyst and electrode screening, bridging the gap between computational and experimental data.
Main
Molecular adsorption energy has been widely used as a descriptor for high-throughput screening of heterogeneous catalysts and electrode materials. [1, 2, 3, 4, 5, 6, 7, 8] Due to the substantial inherent uncertainties arising from the facts, such as side reactions, surface defects, and impurities in most experiments, precise experimental adsorption energies could only be obtained through single-crystal adsorption microcalorimetry with an elaborate experimental design so far, resulting in scarce benchmark experimental data. [9, 10, 11, 12, 13, 14, 15, 16, 17] This limits the application in screening candidate materials. Consequently, the adsorption energies derived from density functional theory (DFT) calculations have been extensively applied for materials screening. [18, 19, 20, 21, 22] Yet, it is not easy to afford the huge computational cost of DFT calculations to achieve a great number of adsorption energies at various active site ensembles in multi-component alloys. Therefore, machine learning models, which use a training dataset labelled by DFT calculations, have recently emerged as a promising method for predicting the adsorption energy () in material screening. [23, 24, 25, 26, 27, 28, 29, 30, 31] Furthermore, the deep learning model demonstrated superior generalization and transferability for a broader range of adsorbates and catalysts compared with surrogate models utilizing a small dataset for specific adsorbates and materials. [32, 33, 34, 35, 36] Notably, machine learning interatomic potentials, which obtain optimised adsorbate structures and adsorption energies using structure coordinates as inputs, have been established as a critical technique in recent years [37, 38, 39, 40, 41].
However, the prediction of chemically accurate adsorption energies using labelled DFT calculation data remains a significant challenge. For one thing, the reduced symmetry of surface adsorption structures increases the computational cost of DFT calculations to label each data point compared to bulk materials and molecules. Furthermore, the low symmetry requires more than two orders of magnitude more training data for the model of surface than for bulk materials and molecules. [42] Consequently, this results in scarcer labelled adsorption data compared to molecular and bulk material data, despite the development of several databases such as Open Catalyst and Catalysis-Hub. [43, 44, 45] Furthermore, it is challenging to leverage limited labelled adsorption data for model training. The labelled DFT data frequently originate from different exchange-correlation functionals and calculation parameter settings, such as different basis sets, k-point samplings, and versions of pseudopotentials, indicating inconsistent and discontinuous potential energy surface data for the training of a machine learning interatomic potential. As a result, only a small portion of the labelled data with consistent exchange-correlation functional and calculation parameter settings in the databases was utilised in most of the previous work in a single task. [46, 47, 48, 49, 50]
More importantly, the accuracy of the predicted based on the deep learning model trained by DFT data is subject to the exchange-correlation functional utilised for the training set, rendering it functional-dependent. The constraint of computational cost renders the adsorption databases to be predominantly composed of the results based on generalised gradient approximation (GGA) exchange-correlation functionals up to now, particularly the PBE [51] functional. [52, 53] However, the PBE functional exhibits limitations in terms of accuracy, exemplified by the classic ”CO puzzle” problem. [54, 55, 56, 57, 58] Moreover, the meta-GGA functionals of SCAN [59] and r2SCAN [60], which have been regarded as the promising functionals to establish a reliable adsorption database recently, even overbind slightly compared to PBE and mistakenly predict the preferential site for CO adsorption as well. [61] The substantial computational cost associated with high-ladder quantum chemistry methods, such as random phase approximation (RPA) [62], which can yield chemically accurate adsorption energies using the embedding strategy [63, 64, 65, 66], precludes their use in constructing a sufficiently large dataset for model training. The -machine learning method presents an alternative approach for predicting chemically accurate adsorption energies for specific systems by incorporating RPA calculation results. Nevertheless, approximately one hundred RPA results are still required to train a model to predict the CO adsorption energy ((CO)) on a specific metal surface using the -machine learning method. [67, 68, 69, 70] Consequently, the extension of the method to vast chemical compound spaces is hindered by significant computational expenditures. Thus, the application of this method to catalyst and electrode screening is severely limited due to these constraints. It is imperative to develop a new model and protocol to predict chemically accurate adsorption energies, making full use of the limited GGA-level data available.
To this end, breaking the limitation of functional dependency is a prerequisite for a new machine learning model for chemically precise prediction. The machine learning interatomic potential, which corresponds to the mapping from input atomic coordinates to output energies, must be functional-dependent, arising from the intrinsic functional-independence of atomic coordinates, while the functional-dependency of energies (Fig. 1a). This indicates that a functional-independent deep learning model may be achieved by employing functional-dependent physical quantities as both model input and output. Notably, both adsorption energies and electronic structure characteristics exhibit strong functional dependency. The density of state (DOS) serves as an information-dense descriptor that compresses a substantial amount of spatially distributed electronic structure information into a computationally affordable one-dimensional energy space, making it a promising candidate as a model input to overcome the limitation of functional dependency. Interestingly, DOS is one of the most frequently used physical quantities that helps to understand the chemisorption strength in heterogeneous catalysis. In particular, the d-band centre theory [71, 72], which is derived from the approximation of the Newns-Anderson model [73, 74, 75], is prevalent in the field of heterogeneous catalysis. Recently, Xin’s group proposed a Bayesian learning method to emphasise the significance of the orbital orthogonalisation term in the Newns-Anderson model for noble metal surfaces, which was previously omitted in the d-band centre theory. [76] Fermi softness offers an alternative explanation for the orbital contribution to adsorption by extending the frontier molecular orbital theory to the solid surface. [77] This approach underscores the critical role of the DOS near the Fermi levels. Fung et al. proposed a deep convolutional neural network (DOSNet) to predict the at the GGA level by utilising the local density of states (LDOS) of the adsorbate and the binding sites of the substrate, mainly based on the adsorbed structures. [78] Nevertheless, the CNN featurizer in DOSNet suffers from low explainability and long-range interaction neglect [79, 80, 81, 82, 83], which limits the comprehensive understanding of the contribution of different energy levels to the adsorption.
To tackle the fundamental challenge of functional dependency for adsorption strength, herein, we present a new deep learning model, called DOS Transformer for Adsorption (DOTA), to efficiently predict accurate adsorption energies across metallic and intermetallic surfaces. The DOTA architecture synergistically integrates the interpretable multi-head self-attention mechanism of the Transformers framework with LDOS feature engineering to first capture the underlying orbital interaction patterns through pretraining using PBE-level LDOS profiles and adsorption energies without prior physical constraints. The functional-independent orbital interaction modes enable transfer learning across quantum chemistry methods and experimental data. Therefore, the chemically accurate adsorption energies of a specific adsorbate across various metallic and intermetallic surfaces could be predicted by fine-tuning the PBE-level model through incorporating the LDOS profile of the gaseous adsorbate at the HSE06 [84] level and less than five precise experimental adsorption energies. In addition, we comprehensively revisited the d-band centre theory and Fermi softness through our interpretable attention paradigm and deciphered the CO/Pt(111) puzzle.
Results
Workflow of DOTA
To enable efficient catalyst screening without requiring preliminary time-consuming geometry optimisation during the inference, DOTA strategically processes gaseous adsorbate LDOS and bare surface LDOS instead of adsorbed complex configurations. The model architecture implements quantum mechanically informed feature engineering, allocating 32 angular momentum-projected DOS (PDOS) embedding channels per surface atom and 8 PDOS embedding channels per adsorbate atom. For instance, the bridge-site adsorption configuration of CO on Pt(111) corresponds to 64 channels for the surface two Pt atoms (2 surface atoms × 32 channels) and 8 channels for the adsorbate atom (C) involved in the bond formation. This channel design ensures compatibility with f-electron systems and spin-polarised calculations. Initial feature processing involved an average pooling layer to normalise energy interval disparities across different DOS calculations. Surface and adsorbate channels were then projected into separate latent spaces through weight-shared fully connected layers with nonlinear activation functions. The multi-head attention mechanism in the Transformer architecture enables efficient modelling of long-range orbital interactions across the entire energy spectrum, capturing the interactions between orbitals with different angular momentum quantum numbers. Final energy predictions are generated by merging surface atom features through flattening and sequential fully connected layer processing. Comprehensive implementation details regarding network architecture and hyperparameter selection are provided in the Methods section.
The DOTA framework employs a two-stage training protocol comprising pretraining and fine-tuning steps to predict chemically precise adsorption energies, as schematically illustrated in Fig. 1b. During the pretraining step, all the LDOS inputs and corresponding adsorption energy outputs were derived from PBE results, thereby constructing a DOTA-PBE model. The pretraining dataset contains 23,861 PBE adsorption energies ((PBE)) involving 11 different adsorbates, which encompass monoatomic adsorbates (H, C, N, O, and S) and their hydrogenated counterparts (CH, , , NH, OH, , and SH) across 1,982 distinct metallic and intermetallic (111) surfaces comprising 37 metallic elements. The selected adsorbates cover molecules with varying degrees of bonding unsaturation, enabling the chemisorption energies to span from near 0 eV to over -10 eV. This broad energy range provides a comprehensive sampling of the diversity of adsorption strengths, ensuring adequate coverage of chemical environments relevant to catalytic processes. The common latent orbital interaction patterns between adsorbates and metallic surfaces are expected to be learnt through the pretraining based on these PBE data. Meanwhile, it is expected that the (PBE) across metallic and intermetallic surfaces could be predicted by the DOTA-PBE model.
The predictive model for the accurate adsorption energies of the studied adsorbate was subsequently achieved by fine-tuning the DOTA-PBE model with the LDOS- data pairs of the studied adsorbate. The multi-head fine-tuning strategy was implemented. One dataset is composed of the GGA-level LDOS- data pairs, in which each data pair was evaluated by the PBE, RPBE, or PBEsol functional, respectively. It is expected that the specific orbital interaction patterns for the studied adsorbate could be learnt from these GGA-level data. The other dataset for fine-tuning only includes a tiny number of hybrid data pairs, generally less than 5. Each hybrid data pair consists of the input HSE06-level DOS profile of the gaseous adsorbate molecule (LDOS(g, HSE06)), the input PBE-level LDOS profile of the bare surface (LDOS(surf, PBE)), and the corresponding experimental (exp.) with high fidelity as the output.
Performance of DOTA-PBE
Fig. 2a illustrates the predictive performance of the PBE-level DOTA model via 5-fold cross-validation on the pre-training PBE dataset. The Transformer-based architecture achieves exceptional accuracy of the DOTA-PBE model with a mean absolute error (MAE) of 0.066 eV and a mean absolute percentage error (MAPE) of 1.53 across all the adsorption systems. Adsorbate-specific analysis reveals MAE ranging from 0.042 eV for monoatomic H to 0.096 eV for . MAE differences among different adsorbates arise from the bonding unsaturation, resulting in different adsorption strengths. The stronger the adsorption, the higher the MAE, but the lower the MAPE. This complementary behaviour of error metrics exhibits the necessity of taking both absolute (MAE) and relative (MAPE) error measures into account when assessing model performance, particularly given the wide energy range in the dataset. Notably, the reliance on electronic structure features rather than geometric parameters enables effective transferability between adsorbates with analogous bonding characteristics. We also tested the OH adsorption atop (111) surfaces. The tested surfaces were not included in the previous pretraining PBE dataset. As shown in Fig. 2b, the d-band centre theory fails at the sites composed of the group 13-14 elements that rely on the s and p orbitals for bonding, while the accuracy of the DOTA-PBE model is element-independent across the metallic and intermetallic surfaces (Fig. 2c). Moreover, it indicates that hydrogenated derivatives may be accurately modelled with minimal additional training data. The electronic structure-based approach of DOTA exhibits robust predictive capability across diverse chemical environments.
We further compared the performance of the DOTA-PBE model with the d-band centre and Fermi softness models. We first take the chemisorption of OH over the top sites of the Ag(111) and Au(111) surfaces as an example, which is not included in our pretraining dataset as well. As shown in Fig. 3a, the d-band of the Ag(111) surface is mainly concentrated in the energy window far below the Fermi level, implying weakened adsorption if following Fermi softness theory. This is not in good agreement with the PBE calculation result of -1.67 eV for the adsorption of OH atop the Ag(111) surface. Moreover, the decomposition of the contribution of each energy level (Fig. 3b) by the integrated gradients algorithm suggests that those energy levels far below the Fermi level made the main contribution for OH adsorption atop Ag(111) rather than those energy levels around the Fermi level. Interestingly, the lower d-band centre of Ag(111) ( = -3.61 eV), compared with the Au(111) surface ( = -3.00 eV), instead leads to the stronger adsorption of OH atop Ag(111) than Au(111) (-1.01 eV). These errors may be attributed to the fact that the electron transfer from the metallic surface to the singly occupied molecular orbital of OH [85] or Pauli repulsion [86], rather than the orbital overlapping between OH and the Ag(111) or Au(111) surface, dominates the chemisorption of OH atop Ag(111) or Au(111), which is difficult to be described by the d-band center or the Fermi softness theory. In contrast, our DOTA-PBE model, which omits any predefined restriction, can accurately predict the adsorption energies of OH atop Ag(111) and Au(111) surfaces, with a prediction of -1.71 eV and -1.13 eV, respectively.
The H adsorption over Pt3Y(111) is a classic example that d-band centre theory fails to predict the adsorption site. [77] PBE calculations show that the (H) at the Pt top site is -2.80 eV, while it is -0.81 eV atop the Y site, indicating that H tends to preferentially adsorb at the Pt site. However, the Pt site has a lower d-band centre ( = -0.32 eV) than the Y site ( = 0.23 eV), indicating stronger adsorption at the Y site if following the d-band centre theory. On the contrary, the larger Fermi softness of the Pt site (Sf = 0.92 ) compared with the Y site (Sf = 0.30 ) indicates the stronger adsorption atop the Pt site if following the Fermi softness theory. Hence, the Fermi softness theory can well describe the H adsorption at Pt3Y, while the d-band centre theory fails. The DOTA-PBE model can also accurately predict this case, with the predicted adsorption energies atop the Pt and Y sites being -2.75 eV and -0.88 eV, respectively. This surface is not included in our pretraining PBE dataset as well. The contribution of each energy level to the H adsorption is elucidated by the integrated gradients algorithm. [87] Fig. 3c shows that the Pt valence bands near the Fermi level account for the H chemisorption atop the Pt site, well aligning with the model of Fermi softness, while the conduction bands of Y play an important role in the H chemisorption atop the Y site. Notably, the upshift of the Y conduction band could lead to the upshift of the d-band centre synchronously, thereby predicting a stronger (H) by d-band centre theory. However, the upshift of the Y conduction band would widen the energy gap between the H 1s orbital and the Y conduction band, constraining their orbital overlap. Consequently, the higher the d-band centre, the weaker the H adsorption atop the Y site, which contradicts the d-band centre theory. It indicates that the d-band centre model is not suitable for describing the system, where the conduction band serves a main role in . Hence, the DOTA-PBE model exhibits better generalisation for the prediction of adsorption energies across metallic and intermetallic surfaces compared with traditional d-band centre theory and Fermi softness theory.
Predictive chemically accurate adsorption energy
Building upon the robust performance and transferability of DOTA-PBE, we further fine-tuned the model employing the multi-head strategy to predict chemically accurate adsorption energies for specific adsorbates across metallic and intermetallic surfaces. The notable scarcity of chemically accurate data requires a model architecture with a minimum requirement of labelled chemically accurate data to fine-tune the model. Taking CO adsorption as an example, we first incorporated 988 LDOS-(CO) pairs (Table S2) across different sites of 174 metallic and intermetallic surfaces, evaluated by PBE, RPBE [88], and PBEsol [89] functionals, respectively, as one fine-tuning dataset. Data derived from different functionals expanded the perceptive chemical space, thereby enhancing the ability of the model to learn orbital interaction patterns between CO and metallic surfaces. Importantly, the foundation of PBE-level orbital interaction understanding, combined with GGA-level insights into the interactions between CO and the surface, enabled exceptional data efficiency to achieve chemical accuracy predictions with only four high-fidelity data points in the other fine-tuning dataset. Each high-fidelity data point comprised experimentally measured CO adsorption energies ((exp.) [18]: Pd(111), Cu(111), Ir(111), or Ni(111)) as output, paired with corresponding input PBE-calculated surface LDOS profiles (LDOS(surf, PBE)) and HSE06-calculated gaseous CO LDOS (LDOS(CO(g), HSE06)) characteristics (Fig. S1a). We compared (CO) over Pt(111) and Rh(111) (Fig. 4a and Fig. 4b) predicted by the DOTA model with different input functional combinations for adsorbate/surface LDOS profiles, the standard PBE, PBEsol, HSE06, RPA, and experimental results. As introduced from Blyholder’s model [90], CO adsorption is primarily formed through the -donation between CO HOMO (5 occupied orbitals) and the metal conduction bands and the d- backbonding between CO LUMO (2 unoccupied orbitals) and the metal valence bands. Accordingly, the correct (CO) and binding sites rely on the correct description of the electronic structures of CO and the metallic surfaces. Fig. 4a shows that both the standard PBE and PBE/PBE combinations results in CO overbonding and mistakenly predicted fcc sites for CO chemisorption over these two surfaces. More poor results were predicted by the standard PBESol and PBESol/PBESol combinations. It was consistent with the previous report [91, 92] that the significantly underestimated HOMO-LUMO gap of CO by PBE and PBEsol functionals [93, 94] should account for the overbinding of CO over the Pt(111) and Rh(111) surfaces. Nevertheless, PBE exhibit good performance at the description of the metallic surface electronic structure, as demonstrated by the closer metal cohesive energy to the experimental value (Fig. 4a and Fig. 4b). The standard HSE06 functional is also poor at predicting CO adsorption because it fails to describe the delocalised electronic structures of metallic surfaces. [95, 84] Interestingly, despite the accurate (CO), neither the CO HOMO-LUMO gap nor the metal cohesive energies could be accurately calculated by RPA (Fig. S1b, Fig. 4a, and Fig. 4b), which is consistent with previously reported results. [58] Importantly, HSE06 exhibits the best predictive performance for CO HOMO-LUMO energy levels, while the PBE functional excels at the prediction of metal cohesive energy [58], indicating that the HSE06 and PBE functionals are good at the description of the molecular orbitals and metal surfaces, respectively. Consequently, the input of LDOS(CO(g), HSE06) and LDOS(surf, PBE) (HSE06/PBE) could provide superior results, with the predicted differing by less than 0.04 eV from the experimental values. More importantly, the model with the HSE06/PBE input combination could correctly predict the top site as the preferential binding site for CO adsorption, aligning with the experimental results. [18] This also manifests that the underestimated HOMO-LUMO gap by the PBE functional is mainly responsible for the ”CO puzzle”, which was speculated by previous reports. [96]. Moreover, the accuracy of the multimodal integration capitalises on the advantages of the experimental precision of microcalorimetry in , the reliability of the PBE functional for the description of electronic structures of metallic surfaces [97], and the superior performance of HSE06 beyond RPA for CO molecular orbital description [95, 84].
Encouraged by the chemically accurate prediction for CO, we further examined the transferability of the fine-tuned DOTA framework to dissociative adsorption of and , under extremely data-limited regimes. For these molecules, the was computed as , where (X = H or O) was predicted by DOTA from atomic LDOS inputs, and or D(O=O) denotes the standard gas-phase bond dissociation energy [98]. Due to the scarcity of high-fidelity experimental dissociative adsorption energies, the model was trained on only 3 data points for H and 2 for O, and validated against 5 independent measurements for each. As shown in Fig. 4, predictions for both training and validation sets cluster closely around the parity line across all three molecules based on minimal supervision. Moreover, the MAE value of 0.033 eV is within the chemical accuracy. These preliminary results suggest that DOTA’s orbital-aware architecture enables robust few-shot learning not only for molecular adsorption but also for dissociative adsorption, highlighting its potential for modelling more surface reactions such as hydrogen evolution or oxygen reduction.
Discussion
Established upon two key physical insights, the orbital-aware adsorption strength and the functional-independent LDOS-adsorption relationship, the DOTA model permits the alignment of experimental sources and multi-fidelity DFT data. It could further achieve chemical accuracy in predicting molecular adsorption energies across extensive metallic and intermetallic surfaces, primarily trained on GGA-level datasets augmented by minimum high-fidelity experimental data. This demonstrates its exceptional transferability among elements. Thus, DOTA represents a transformative leap in computational catalysis by addressing longstanding challenges such as functional dependency, limited availability of high-fidelity data, and the need for interpretable machine learning models. By synergising deep learning architectures and physically informed features like LDOS, it achieves unprecedented accuracy and generalisability across chemical element environments.
Herein, the limited high-fidelity experimental data from the (111) surface render the DOTA model to be trained upon the data from the flat (111) surface. Because the orbital interaction patterns between the adsorbate and metallic surface would vary with the exposed facet, the DOTA model would be sensitive to the surface structures. In addition, the available high-fidelity experimental data correspond to a specific surface coverage, necessitating consistency with the GGA data for fine-tuning. Nevertheless, with the rapid accumulation of GGA data, the capacity of DOTA to unify multi-fidelity data sources enables the rapid construction of a large dataset of adsorption energies with high fidelity over different exposed surfaces, requiring an extremely small quantity of extra high-fidelity experimental or high-ladder quantum chemistry data. This could significantly accelerate the discovery of next-generation catalysts and electrode materials.
Methods
All the DFT calculations were performed with Vienna ab initio Simulation Package (VASP) [99, 100] with periodic boundary and the projector augmented wave (PAW) pseudopotentials [99, 101]. The Catalysis Kit (CatKit) software [44] and the Atomic Simulation Environment (ASE) Python library [102, 103] were used to generate and classify the adsorption structures. The chemisorption structures were obtained in reference [45] for DOTA model pretraining, where the PBE functional was used for both and DOS profile calculations. For the fine-tuning stage, PBE, RPBE and PBEsol functionals were used for and DOS profile calculations. The hybrid HSE06 functional was used only for the DOS profile of the gaseous adsorbate. All calculations were conducted with spin polarisation, using a kinetic energy cutoff of 450 eV. The p(22) 3-layer slab was employed for each metallic and intermetallic surface. The corresponding Monkhorst-Pack k-point grid [104] was 771, with a first-order Methfessel-Paxton smearing of 0.1 eV. The resolution of calculated lm- and site-DOSs with reference to the Fermi level, which were employed for DOTA training, was 0.01 eV.
The DOTA model was implemented with the PyTorch Python library [105]. The AdamW optimiser with an initial learning rate of 0.0002 and the L1 loss were used for training. An exponential cosine annealing scheduler was implemented to adjust the learning rate, with 150 epochs used, comprising 50 epochs for warming up. A batch size of 256 was employed. The contribution of each energy level to the was calculated using the interpretable integrated gradients algorithm implemented in the Captum library [87] with default parameters.
Code availability
Code for DOTA is available as an open-source repository on GitHub, https://github.com/zhzhang321/DOTA
Acknowledgements
This work was supported by the National Key Research and Development Program of China (2023YFA1507601 and 2021YFA1500700). We thank the Centre for High Performance Computing at Shanghai Jiao Tong University for providing the computing resources of the Siyuan-1 cluster.
Author information
Author contributions
Z.Zhang wrote the code, performed theoretical simulations, and wrote the paper. X.-M.C. designed the study, analysed the data, and wrote the paper. Both authors discussed the results and improved the manuscript.
Corresponding authors
Correspondence to Xiao-Ming Cao.
Ethics declarations
Competing Interests
All authors declare no competing interests.
Supplementary information
Supplementary Figs. 1-3 and Tables 1-4.
Rights and permissions
Reprints and permissions information is available at http://www.nature.com/reprints.